OpenStudy (anonymous):

What expression is the simplified form of square root of y^4 A. y B.y^2 C.y square root y D. y^2 square root y

OpenStudy (anonymous):

think of y^4 as yyyy now, how many pairs of y are there?

OpenStudy (anonymous):

\[\sqrt{y^{4}} = \sqrt{yyyy}\]

OpenStudy (anonymous):

you can pull out every pair you find, there are 2 pairs of y in there, so: \[yy \sqrt{1} = y^2\]